OpenStudy (anonymous):

How can logarithms be graphed with different bases? How do transformations affect the logarithmic graph?

OpenStudy (dan815):

do you know the relationship between f(x) and f^-1(x)

OpenStudy (empty):

Well the simplest way to think of different bases is to exploit this: \[y=\log_b(x) = \frac{\log_a(x)}{\log_a(b)}\] So if you want to compare the bases a and b then you can always write one in terms of the other scaled by a factor.

OpenStudy (anonymous):

No

OpenStudy (dan815):

graphing y=f(x) and y=f^-1(x) is same as graphginh y=f(x) and x=f(y)

OpenStudy (dan815):

you just look at your axis flipped

OpenStudy (dan815):

and log is the inverse of exponential function

OpenStudy (anonymous):

@Empty

OpenStudy (dan815):